The length of a rectangle is 3 times the width. The perimeter is 96 cm. Find the width and length.
Answer by checkley71(8403)

PERIMETER IS EQUAL TO 2 TIMES THE LENGTH PLUS 2 TIMES THE WIDTH
THUS P=2L+2W OR 96=2*3W+2*W OR 96=6W+2W OR 96=8W OR W=96/8 OR W=12 THUS
86=2*L+2*12 OR 96=2*L+24 OR 96-24=2L OR 72=2L OR 72/2=L OR L=36
PROOF 86=2*36+2*12 OR 86=72+24 OR 86=86 WHEN THE LENGTH (36) IS 3 TIMES THE WIDTH (12) 3*12=36.